OpenStudy (anonymous):

Find all solutions in the interval [0, 2π). sin2 x + sin x = 0

OpenStudy (anonymous):

@SolomonZelman ?

OpenStudy (michele_laino):

please use this identity: \[\sin(2x)=2 \sin x \cos x\]

OpenStudy (anonymous):

so 2sinxcosx+sinx =0 ?

OpenStudy (solomonzelman):

no exactly, you forgot the `sin x` that you have in your equation already.

OpenStudy (michele_laino):

ok! now factorise, putting in evidence sin (x), please

OpenStudy (anonymous):

im not sure what that means

OpenStudy (michele_laino):

I write: \[\sin x(2\cos x+1)=0\]

OpenStudy (solomonzelman):

here is what you did. \(\large\color{black}{ \color{red}{ \sin(2x)}- \sin x = 0}\) \(\large\color{black}{ \color{red}{ 2\sin(x)\cos(x)}~~~~~ = 0}\) but you left something out, didn't you?

OpenStudy (anonymous):

okay i see, now what?

OpenStudy (solomonzelman):

first, write the equation that you are up to correctly, please.

OpenStudy (anonymous):

you need the +sinx after the cos(x) @SolomonZelman

OpenStudy (michele_laino):

now, please you have to apply the Law of cancellation of the product

OpenStudy (michele_laino):

so you got two equations, namely: sin x=0, and 2 cos x +1=0

OpenStudy (solomonzelman):

ohh it is plus, not minus. so yes and you would be getting, \(\large\color{black}{ \color{red}{ \sin(2x)}+ \sin x = 0}\) \(\large\color{black}{ \color{red}{ 2\sin(x)\cos(x)}+\sin(x)= 0}\)

OpenStudy (solomonzelman):

then, factor out of sin(x), since it is the common factor in both terms.

OpenStudy (michele_laino):

\[\sin x=0, 2 \cos x+1=0\]

OpenStudy (anonymous):

so cosx= 1/2?

OpenStudy (michele_laino):

no, cos x=-1/2, because you have to add -1 to both sides of the second equation first

OpenStudy (solomonzelman):

\(\large\color{black}{ 2\color{blue}{\sin(x)}\cos(x)+\color{blue}{\sin(x)}= 0}\) \(\large\color{black}{ \color{blue}{\sin(x)}(2\cos(x)+1)= 0}\)

OpenStudy (solomonzelman):

So, either \(\large\color{black}{ \color{blue}{\sin(x)} =0}\) or \(\large\color{black}{ 2\cos(x)+1= 0}\)

OpenStudy (anonymous):

okay so then do i look at the unit circle or is there another step?

OpenStudy (solomonzelman):

well, for the second option you need a cos(x) that is equal to -1/2...

OpenStudy (michele_laino):

please, there are no steps

OpenStudy (anonymous):

so is it 0, 2pi/3, 4pi/3?

OpenStudy (michele_laino):

that's right!, there is another solution, namely x= pi, which derives from sin (x)=0

OpenStudy (anonymous):

now i am confused because that is not one of the answer choices? let me list the choices really quick

OpenStudy (anonymous):

0, pi, pi/3, 5pi/3 0, pi, pi/3, 2pi/3 0, pi, 4pi/3, 5pi/3 0, pi, 3pi/2

OpenStudy (michele_laino):

please note that: x=0, pi are the solution of sin (x)=0, whereas, x=2pi/3, 4pi/3 are the solutions of 2 cos(x)+1=0

OpenStudy (michele_laino):

sorry , x=2pi/3, 5pi/3 are the solutions to the equation 2 cos(x)+1=0

OpenStudy (anonymous):

so what do i do?

OpenStudy (michele_laino):

sorry it I wrote before, namely solutiuons are: x= 0, pi, 2pi/3, 4pi/3

OpenStudy (anonymous):

but those arent an answer i can choose

OpenStudy (michele_laino):

I see, nevertheless those solutions are right

OpenStudy (anonymous):

i know but i dont know how to answer it because its an online class so i cant really fight that the answer choices are incorrect...

OpenStudy (michele_laino):

you can check the correctness by substituting those solutions in your equation

OpenStudy (anonymous):

how do i type in the sin^2 part

OpenStudy (michele_laino):

why sin^2?

OpenStudy (anonymous):

okay when i plugged it in it says the only ones that equal 0 are the last choices

OpenStudy (michele_laino):

sorry is your equation this? \[(\sin x)^{2}+\sin x=0\]

OpenStudy (anonymous):

its sin^2 x

OpenStudy (anonymous):

is there a way you can help me with another question?

OpenStudy (michele_laino):

thn yur equation is that I wrote above, yes!

OpenStudy (anonymous):

ill post it in another question and tag you

OpenStudy (michele_laino):

you ave to factor oout sin(x), so your equation will ecome: \[\sin x(\sin x +1)=0\]

OpenStudy (michele_laino):

so solution s are: x=0, pi fro equation sin x=0, whereas x0 3pi/2 from equation sin x+1=0 so the last option is your answer

OpenStudy (michele_laino):

@adekker